# finantial support acknowledgement

### v2

@@ III-C2 Targeted Backdoor Attack
This attack is based on [3]. Each malicious client owns one mislabeled image samples and trains the local model on it. We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that side-job training (i.e., backdoor training) remains stealthy. Then, the attack boosted malicious clients’ updates to negate the combined effect of the benign client:
$w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$
where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor. In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ’airplanes’ as ’bird’ whereas normally classifying the image of other classes like ’cat,’ ’horse,’ or ’truck.’
IV FedCC: Robust Aggregation Against Poisoning Attacks
Figure 2: An Overall Mechanism of our FedCC.
IV-A Overview of FedCC
@@
Figure 2 depicts an overview of FedCC. (1) A server initializes and advertises a global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$. (2) The clients individually train their local models using their respective datasets and send the trained weights to the server for aggregation. Our proposed method intervenes at this moment. (3) The server extracts PLRs from the global model and each local model, and (4) compares RBF Kernel CKA values between them. (5) The CKA values are then clustered into two, potentially representing benign and malicious clients. (6) Since the server cannot assure whether the clients with higher similarities are malicious, it performs a coordinate-wise average of the weights from clients belonging to the cluster with a larger number of members, backed by an assumption that the number of malicious clients does not exceed $n/2-1$. (7) Finally, the server updates the global model for the current epoch and proceeds to the next iteration by distributing the updated global model to newly selected $n$ clients.
IV-B Detailed Design
@@
When filtering malicious clients in the presence of non-IID data, a major challenge is the server’s lack of knowledge regarding the data distribution. Specifically, even if the data is benign, local models can exhibit significant angular or magnitude differences if the data is non-IID. We thus require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
@@
TABLE I: Comparison of Performance with Various Similarity Metrics
@@
Penultimate Layer is the output of the second-to-last layer before the softmax layer in CNN. In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients. Therefore, the collected weights are the only available information for determining whether a client is malicious. PLR takes the end result of the convolution and pooling process that will reach a classification decision. Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
Kernel CKA[13] is a similarity metric to assess the similarity between two neural networks. It compares representations at different layers or models trained in various ways. The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
To justify the use of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine (angle), and Euclidean similarity (distance). We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings. The results are summarized in Table I. In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance. We observed that Kernel CKA consistently yielded the highest performance across multiple experiments. Therefore, we decided to utilize Kernel CKA to compare PLR similarities.
Algorithm 1 FedCC
Input global_w, n local_w $w_{1}$, …, $w_{n}$
      Output agg_w
1:for $i<n$ do $\triangleright$ Get PLRs of local and global models
2:     local_plrs[$i$] $\leftarrow$ local_ws[second_last_layer]
3:end for
4:glob_plr $\leftarrow$ global_w[second_last_layer]
5:for $i<n$ do $\triangleright$ Apply kernel CKA to the plrs
6:     cka[$i$] $\leftarrow$ cka(glob_plr, local_plrs[$i$])
7:end $\triangleright$ Apply Kmeans clustering algorithm
8:kmeans = kmeans(n_cluster=2, cka[$i$])
9:labels = kmeans.labels_
10:count = counter(labels)
11:suspect $\leftarrow$ 1
12:if count[00] $\leq$ count[$1$then
13:     suspect $\leftarrow$ 0
14:end $\triangleright$ Define suspect based on the size of cluster
15:suspects = where(label == suspect)
16:for $i$ in range($n$do
17:     for $s$ in suspects do $\triangleright$ Set scale to 0 for the suspects
18:         scale[$s$] $\leftarrow 0$
19:     end for
20:     if scale[$i$] != 0 then
21:         scale[$i$] $\leftarrow 1$
22:         selected_param.append(scale[$i$] * local_ws)
23:     end if
24:end $\triangleright$ Average the selected parameters
25:agg_weights = mean(scale*selected_param)
26:return agg_weights
@@
FedCC is an aggregation method that combines selected weights without scaling, based on the measurement of Kernel CKA similarities between the global and local models’ PLRs. A pseudo algorithm is described in Algorithm 1. A server first extracts PLRs of the global and local models and calculates the kernel CKA between them. We specifically focus on comparing PLRs because they exhibit the most distinguishable differences among differently trained models, as demonstrated in previous sections. Additionally, we utilize RBF Kernel CKA, as it effectively captures similarity differences between non-IID and malicious clients. The similarity values obtained are then clustered using a simple K-means clustering algorithm. Any clustering method that produces only two clusters can be used. Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the cluster with fewer members as suspects. The indices of these suspicious clients are stored in a list called ‘suspects.’ We proceed by selectively averaging the weights of clients whose index does not belong to the ‘suspects’ list, excluding the weights of the suspicious clients from the aggregation. Note that the weights are averaged without scaling, as scaling could impede the model from convergence even when the loss reaches optima. We named our proposed method FedCC.
Compared to other methods such as Krum, Bulyan, and FLARE, which either require prior knowledge of the number of attackers or allow the server to have access to raw data, FedCC does not rely on any information about the clients or any instances of data. This ensures strong privacy preservation within the federated learning framework. Furthermore, FedCC effectively distinguishes between malicious and benign clients, even in practical but challenging scenarios where benign clients have non-IID data.
V Experiments
V-A Dataset
V-B Non-IID Simulation
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we use Dirichlet distribution, which is well-known for its capability of representing real-world data [15]. A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class.
V-C Model Architecture
@@
Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN architectures for the experiments. For fMNIST, we used a 4-layer-CNN with two convolutions and two fully connected (FC) layers, incorporating dropout with probabilities of 0.25 and 0.5 before and after the FC layer, respectively. For CIFAR10, we utilized a 5-layer-CNN with three convolutions, each followed by a max-pool layer and two FC layers, applying dropout with a probability of 0.5 before the first FC layer. Finally, for CIFAR100, we adopted the LeNet architecture. Table VI, Table VII, and Table VIII in Appendix summarize model architectures for each benchmark dataset.
V-D Setup
@@
We implement two untargeted attacks in [8] and one targeted attack [3], defense baselines, including Krum [4], Coomed [26], Bulyan [11], and FLARE [22], and FedCC. We have ten clients with a client participation fraction of 1.0 unless stated otherwise. The number of attackers is 30% of the total clients (three) in untargeted and 10% (one) in targeted attacks as in [22]. We also assumed all clients have the same NN architecture. Under the targeted backdoor attack, benign clients go through three, whereas compromised clients undergo six local epochs. We train each client’s local model using an Adam optimizer with a learning rate of 0.001, following the experiment in [22]. All results are averaged after running each experiment three times.
TABLE II: Accuracy of no-attack scenario, boldface and underlining refer to the highest and the second highest accuracy, respectively.
Scenarios Dataset FedAvg Krum Coomed Bulyan FLARE FedCC
No Attack non-IID fMNIST 69.68 43.10 76.12 68.78 42.82 71.92
CIFAR10 53.06 16.16 45.55 46.67 27.33 50.80
CIFAR100 17.22 7.27 12.82 15.21 0.99 17.62
No Attack IID fMNIST 89.89 83.81 88.91 89.49 76.86 89.71
CIFAR10 69.17 56.56 67.34 67.81 46.84 69.03
CIFAR100 21.49 7.73 15.56 20.55 7.59 22.05
V-E Evaluation Metrics
@@
We assessed the performance of our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy). The confidence indicates the likelihood of misclassification, with values ranging from 0 to 1 (lower values indicating better defense performance). Accuracy measures the overall test accuracy of the global model when various defense methods are applied when attacks exist, with values ranging from 0 to 100 (higher values indicating better performance). We reported both metrics in the presence of backdoor attacks while focusing solely on accuracy for untargeted attacks—as the attacker’s goal is to deteriorate the overall accuracy.
VI Results and Discussion
In Table II, which represents the baseline experiments without any attacks, we can observe the accuracy of different methods in both non-IID and IID settings. FedCC achieves accuracy comparable to that of the best-performing method, FedAvg. The slightly higher accuracy of FedAvg than FedCC may be attributed to its wholesome averaging, which leads to a more generalized model when there is no attack. Nonetheless, FedCC surpasses all other baseline methods with CIFAR100, which has more classes, making it harder to achieve better accuracy in FL[7].
TABLE III: Accuracy under untargeted and targeted attacks in Non-IID (top) and IID setting (bottom)
@@ VI Results and Discussion
Scenarios Dataset FedAvg Krum Coomed Bulyan FLARE FedCC
Fang-Krum non-IID fMNIST 64.63 16.51 57.12 13.39 49.54 66.27
CIFAR10 33.69 15.38 35.70 19.23 17.03 37.40
CIFAR100 2.27 1.00 4.85 0.98 7.46 14.51
Fang-Med non-IID fMNIST 16.32 49.33 68.90 64.12 52.25 72.76
CIFAR10 10.02 25.06 40.23 32.47 14.59 42.37
CIFAR100 1.00 6.24 10.27 6.91 1.00 16.12
Targeted non-IID fMNIST 79.07 43.46 71.25 73.58 60.57 84.54
CIFAR10 30.28 16.05 28.87 25.77 10.52 36.20
CIFAR100 9.84 7.24 11.02 12.01 1.04 13.81
Fang-Krum IID fMNIST 75.55 31.66 87.62 50.30 79.16 89.23
CIFAR10 49.67 40.86 57.40 12.67 25.77 67.12
CIFAR100 13.72 1.04 6.17 1.59 7.49 18.47
Fang-Med IID fMNIST 20.86 85.33 86.70 87.45 71.08 88.44
CIFAR10 9.51 54.28 59.20 57.69 49.82 64.74
CIFAR100 0.87 12.27 14.43 12.56 5.83 17.83
Targeted IID fMNIST 89.32 85.52 88.45 89.35 67.65 90.16
CIFAR10 52.20 45.37 49.27 55.77 10.09 51.57
CIFAR100 12.73 8.37 15.91 16.59 1.71 18.08
VI-A Non-IID Data Environment
Following the aforementioned non-IID simulation, we set $\alpha$ to 0.2 to create a disjoint client training data and distribute each set of data to each client.
VI-A1 Untargeted Model Poisoning Attacks
@@
The upper section of Table III provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) and Targeted attacks in a non-IID environment. One notable finding is that robust aggregation algorithms often yield lower accuracy compared to simple averaging (no defense). This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with benign clients. When applying Krum and Bulyan against Fang-Krum, the accuracy drops by a substantial 59.30% and 65.68%, respectively, compared to FedAvg. This decline can be attributed to Fang-Krum’s intentional design to degrade accuracy against Krum. Additionally, Coomed struggles to effectively mitigate this attack as the median values of weights often fail to represent benign clients, particularly in non-IID settings, accurately. Furthermore, the relatively low accuracy of FLARE is due to weight scaling, even when the model already exhibits satisfactory performance. In contrast, FedCC, which selectively averages weights, stands out by neither boosting nor minimizing weights, resulting in higher performance than other methods. Notably, FedCC achieves the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively), with a significant leap in experiments with CIFAR100.
@@
Turning our attention to the Fang-Med attack, the accuracy of FedAvg experiences a substantial drop due to the attack’s tendency to deviate from median values, creating outliers. This behavior explains the relatively high accuracy when Coomed and Bulyan are applied because coordinate-wise median and trimmed mean methods disregard outliers. Meanwhile, FLARE performs poorly for a reason similar to the one mentioned above. For both attacks, we can observe that Krum, which selects a single local model as the global model, is less robust than Coomed, which aggregates multiple local models. Similarly, Bulyan, which takes averages of multiple chosen clients, exhibits less robustness than Coomed, particularly in non-IID data environments. Notably, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data with the server or other clients with non-IID data. This exceptional performance of FedCC can be attributed to two factors: 1) highly distinguishable information within PLRs, and 2) higher CKA similarity between benign clients than the one between benign and malicious clients.
VI-A2 Targeted Backdoor Attacks
Figure 4: Confidence of Backdoor Task in a Non-IID Setting (top) and an IID setting (bottom)
The fourth row of Table III summarizes accuracy under targeted backdoor attacks in a non-IID setting. To further illustrate the impact of these attacks, we include a visual representation of the backdoor confidence in the first row of Figure 4.
It becomes evident that Krum, which selects a single client’s weights as the global model, exhibits diminished robustness compared to other methods that aggregate the weights of multiple clients; coordinate-wise median (Coomed) and trimmed mean with multi-Krum (Bulyan) demonstrate relatively higher test accuracy. Conversely, FLARE decreases the main task test accuracy for similar reasons observed in the experiments mentioned above. Nonetheless, FedCC achieves the highest test accuracy. This outcome underscores the superiority of filtering malicious clients through CKA similarity comparison, surpassing the effectiveness of first-order statistical methods like mean or median.

### v3

@@ 3.3.2 Targeted Backdoor Attacks
are based on [3]. Each malicious client owns one sample of mislabeled images and trains the local model on it. We trained the model on backdoor tasks along with the main task(s); the training went on for both malicious and benign tasks to maintain benign accuracy such that backdoor training remains stealthy. Then we boosted malicious clients’ updates to negate the combined effect of the benign agent: $w_{i}^{t}=w_{G}^{t-1}+\alpha_{m}(w_{i}^{t-1}-w_{G}^{t-1})\text{ for }i=1,...,m$ where $t$ is the current epoch and $\alpha_{m}$ is a boosting factor. For DBA [33], we implant split backdoor triggers into the input so that the global model can be attacked in a distributed manner. For more detail, refer to [33]. In CIFAR10, for example, we manipulate the model parameters to misclassify an image of ‘airplanes’ as ‘bird’ whereas normally classifying the image of other classes.
4 FedCC: Robust Aggregation against Poisoning Attacks
4.1 Overview of FedCC
@@
Figure 1 depicts an overview of FedCC. (1) A server initializes and advertises the global model to $n$ participating clients, selected out of total $k$ clients with a probability of fraction $C$. (2) The clients individually train their local models using their respective datasets and send the trained weights to the server for aggregation. Our proposed method intervenes at this moment. (3) The server extracts PLRs from both the global model and each local model, and (4) compares RBF Kernel CKA values between them. (5) The CKA values are then clustered into two, potentially representing benign and malicious clients. (6) The CKA similarity values are then normalized across all clients for every layer as $a$, but the second last layer’s CKA similarity values are normalized within the larger cluster as $b$, backed by the assumption that the number of malicious clients does not exceed $n/2-1$. (7) Finally, the server aggregates each layer of local models weighted by $a$, but the PLR weighted by $b$, to update a global model for the current epoch and proceeds to the next epoch by distributing the updated global model to the newly selected $n$ clients.
Figure 1: An Overview of FedCC. $k$, $n$, $a$, and $b$ indicate the kth layer, the number of participating clients, normalized similarity across all layers (i.e., across_cka), and normalized similarity within the cluster (i.e., within_cka), respectively.
4.2 Detailed Design
@@
A major challenge when filtering malicious clients in the presence of non-IID data is the server’s lack of knowledge regarding the data distribution. Even if the data is benign, local learning models can exhibit significant angular or magnitude differences if they are non-IID. Thus, we require a similarity metric independent of data distribution and not influenced by the distance or direction of model parameters.
4.2.1 Penultimate Layer
is the output of the second-to-last layer before the softmax layer in CNN. In federated learning, where the server cannot access local training data, it receives trained model parameters (weights) from selected clients. Therefore, the collected weights are the only available information for determining whether a client is malicious. PLR takes the end result of the convolution and pooling process that will reach a classification decision. Thus, we assumed that computed weights of the second last layer might contain the essential information that strongly influences classification results.
@@ 4.2.1 Penultimate Layer
4.2.2 Kernel CKA[13]
is a similarity metric to assess the similarity or dissimilarity between two neural networks. It compares representations at different layers or models trained in different ways. The sophisticated similarity comparison ability of Kernel CKA makes it a suitable choice for our purpose.
To justify the effectiveness of Kernel CKA, we compared its performance with other similarity metrics such as linear CKA, MMD (not normalized CKA), cosine similarity (angle), and Euclidean similarity (distance). We evaluated the test accuracy of FedCC using the CIFAR10 dataset under two untargeted attacks and one targeted attack, separately in IID and non-IID settings. The results are summarized in Table 1. In the ’targeted’ column, the values represent the combination of test accuracy and backdoor confidence, with higher test accuracy and lower backdoor confidence indicating better performance. We observed that Kernel CKA consistently yields the highest performance across experiments; therefore, we decided to utilize Kernel CKA to compare the similarities of PLRs.
Table 1: Comparison of Performance with Various Similarity Metrics
@@ 4.2.2 Kernel CKA[13]
4.2.3 FedCC
@@
is an aggregation method that combines local clients’ model parameters, based on the Kernel CKA similarities between the global and local models’ PLRs. An overall algorithm is described in Appendix 0.B. A server first extracts PLRs of the global and local models and calculates the kernel CKA between them. We specifically focus on comparing PLRs because they exhibit the most distinguishable differences among differently trained models, as demonstrated in previous sections. Additionally, we utilize RBF Kernel CKA, as it effectively captures similarity differences between non-IID and malicious clients. The similarity values obtained are then clustered using a simple K-means clustering algorithm. Any clustering method that produces only two clusters can be used. Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the larger cluster as ‘candidates.’
Since the earlier layers in CNNs learn global features and the latter layers embed local features [15, 25, 23], we used two differently normalized CKA similarity values: one normalized within the candidates (i.e., within_cka, denoted as $b$ in Figure 1) and one normalized across all the clients (i.e., across_cka, denoted as $a$ in Figure 1). For the PLR (second-last layer), we apply within_cka to the clients belonging to the larger cluster, whereas for all other layers, we apply across_cka to weighted average model parameters. By doing so, we can ensure that the end magnitude of the aggregated parameters stays similar to before the manipulations, leveraging the fact that the sum of the normalized values is one. Moreover, since our method evaluates clients in every global round and does not exclude any local models, it is also effective for adaptive attacks like DBA.
5 Experiments
5.1 Dataset and Model Architecture
@@
We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10 and CIFAR100 [14]. Details about datasets can be found in Appendix 0.A. Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN architectures for the experiments. For fMNIST, we used a 4-layer-CNN with two convolutions and two fully connected (FC) layers, incorporating dropout with probabilities of 0.25 and 0.5 before and after the FC layer, respectively. For CIFAR10, we utilized a 5-layer-CNN with three convolutions, each followed by a max-pool layer and two FC layers, applying dropout with a probability of 0.5 before the first FC layer. For CIFAR100, we adopted LeNet [16].
5.2 Non-IID Simulation
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we use Dirichlet distribution, which is widely used and well-known for its capability of representing real-world data [18]. A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class. We set the $\alpha$ to 0.2 to represent non-IID, based on [11, 38].
5.3 Experimental Setup and Baselines
@@
We implement two untargeted attacks from [7] and two targeted attacks from [3] and [33], along with defense baselines including Krum, multi-Krum [4], Coomed [34], Bulyan [10], FLTrust [6], FLARE [30], and FedCC. We use ten clients with a 1.0 client participation fraction unless stated otherwise. As in [30], we have three adversaries for each untargeted attack and one for the targeted attack. In the DBA, we randomly select four malicious clients and implant five distributed triggers per batch. To compensate for delayed attack effects, we use FedAvg until epoch 30, after which each defense method is applied. Note that, for a fair comparison, FLTrust is computed using its previous global model without auxiliary data. Benign clients undergo three local epochs, while compromised clients undergo six, with training done using the Adam optimizer (learning rate = 0.001).
5.4 Evaluation Metrics
@@
We evaluated our defense mechanism using two metrics: backdoor confidence (confidence) and global model test accuracy (accuracy). The confidence metric measures misclassification likelihood, ranging from 0 to 1 (lower values are better). Accuracy reflects the global model’s overall test performance under defense methods during attacks, ranging from 0 to 100 (higher values are better). Both metrics were reported under backdoor attacks, with a focus on accuracy for untargeted attacks, as the attacker’s goal is to reduce overall accuracy.
6 Results and Discussion
In this section, we report the performance of our FedCC against two untargeted and two targeted attacks in both IID and non-IID environments. We additionally demonstrate the defense effect against a distributed adaptive attack, DBA [33]. Unlike centralized backdoor attacks, the DBA splits a trigger pattern across multiple adversaries, evading the detection by anomaly detection. We show the result of CIFAR10, with various defense methods applied.
6.1 Non-IID Data Environment
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we employ Dirichlet distribution that 0 indicates the most heterogeneity, and 100 mimics homogeneity [20]. A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class. Since $\alpha$ being 0.2 represents a highly non-IID scenario based on [11, 38], we set it as such, following [38].
Table 2: Test Accuracy under untargeted attacks in Non-IID setting.
@@ 6.1 Non-IID Data Environment
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum non-IID fM 64.63 16.51 45.91 57.12 13.39 49.37 49.54 66.27
C10 33.69 15.38 20.09 35.70 19.23 36.63 17.03 37.40
C100 2.27 1.00 5.34 4.85 0.98 11.62 7.46 14.51
Fang -Med non-IID fM 16.32 49.33 66.30 68.90 64.12 12.38 52.25 72.76
C10 10.02 25.06 40.52 40.23 32.47 12.39 14.59 42.37
C100 1.00 6.24 13.97 10.27 6.91 1.00 1.00 16.12
6.1.1 Untargeted Model Poisoning Attacks in non-IID setting
@@
Table 2 provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) in a non-IID environment. A notable finding is that robust aggregation algorithms often yield lower accuracy than simple averaging (fedAvg). This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with those of benign clients. Under Fang-Krum, Krum, Bulyan, and multi-Krum experience significant accuracy drops (16.51%, 13.39%, and 45.91%, respectively, compared to FedAvg’s 64.63%), reflecting the attack’s design to degrade Krum-based defenses. Coomed also struggles, as median weights fail to represent benign clients effectively in non-IID settings. FLTrust mitigates Fang-Krum but performs better on CIFAR10 than fMNIST, likely due to fMNIST’s simpler gradients and lower variance, which challenge its ability to distinguish malicious updates. FLARE’s lower accuracy stems from its approach of scaling the entire model weights uniformly, which ignores knowledge contributed by individual clients. In contrast, FedCC, which selectively averages weights layer-wise, achieves superior performance by boosting or minimizing weights based on their alignment with benign updates. FedCC demonstrates the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively) and shows significant improvement in CIFAR100 experiments.
@@
Turning to the Fang-Med attack, FedAvg experiences a significant drop in accuracy due to the attack’s creation of outliers by deviating from median values. Consequently, Coomed and Bulyan perform relatively well since their coordinate-wise median and trimmed mean methods disregard outliers effectively. FLARE continues to perform poorly for a similar reason mentioned earlier. Between the Krum-based methods, multi-Krum and Coomed are more robust than Krum, as they aggregate multiple local models, providing better resilience to outliers. Similarly, Bulyan’s averaging of multiple clients enhances robustness but falls short of Coomed, particularly in non-IID environments. Finally, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data, showcasing its exceptional performance. This success is attributed to two factors: (1) highly distinguishable information within PLRs and (2) higher CKA similarity between benign clients than between benign and malicious clients.
6.1.2 Targeted Backdoor Attacks
Table 3: Test Accuracy under targeted attacks in Non-IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target non-IID fM 75.65 45.27 65.97 71.70 57.96 61.82 64.31 75.66
C10 36.16 14.98 30.72 48.97 40.11 44.06 10.18 51.56
C100 4.46 6.18 6.90 12.04 11.16 12.95 1.14 15.26
DBA C10 38.56 24.94 7.09 44.45 34.19 51.49 38.73 52.28
Figure 3: Confidence of Backdoor Task for targeted attacks in Non-IID settings.
Table 3 summarizes test accuracy under targeted attacks in a non-IID setting. To further illustrate the impact, Figure 3 provides a visual representation of backdoor confidence. It is evident that Krum, which selects a single client’s weights as the global model, shows reduced robustness compared to methods that aggregate weights from multiple clients. In contrast, Coomed, multi-Krum, and Bulyan exhibit higher test accuracy. FLARE reduces main task accuracy for reasons similar to those observed in previous experiments. FedCC achieves the highest test accuracy. Since targeted attacks aim to misclassify a specific target class while maintaining the main task accuracy, the best performance of FedCC is attributed to leveraging knowledge from earlier stages of training. This result highlights the superiority of using sophisticated weighted knowledge and filtering malicious clients via CKA similarity, which outperforms first-order statistical methods like mean or median. A similar trend is observed with the DBA attack, except for multi-Krum; the lowest accuracy occurs because all four malicious clients, each with a distributed trigger, are treated as outliers and excluded from aggregation.